which equation of the line has a slope of 5 and passes through point (2,1)?
Answer by checkley71(8403)

Y=mX+b m=5 THE SLOPE
Y=5X+b NOW SUBSTITUTE THE GIVEN POINTS FOR Y & X WE GET
1=5*2+b
b=1-10
b=-9
THEN THE EQUATION FOR THIS LINE IS
Y=5X-9